OpenStudy (anonymous):

An SRS of 20 recent birth records at the local hospital were selected. In the sample, the average birth weight was 121.4 ounces and the standard deviation was 7.5 ounces. Assume that in the population of all babies born in this hospital, the birth weights follow a Normal distribution, with some mean μ. The standard error of the mean is A. 27.1. B. 6.1. C. 1.7. D. 0.4.

OpenStudy (kirbykirby):

\[\frac{\sigma}{\sqrt{n}}=\frac{7.5}{\sqrt{20}}\approx1.67\approx.1.7\]